# Finite-sample Identification of Continuous-time Parameter-linear Systems

### v1

@@ 3.1 Building blocks of the filter
As $\left\langle f,g\right\rangle_{N}$ converges to a Riemann integral as $N\to\infty$, so $p_{N}^{d}$ converges too. For later convenience, define the filter-specific constants
@@
$g_{N}^{d}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)\left|p_{N}^{d}\left(\frac{j}{N}\right)\right|$
$=\left\langle p_{N}^{d},\operatorname*{sign}p_{N}^{d}\right\rangle_{N},$
where $\operatorname*{sign}p_{N}^{d}$ is an abuse of notation referring to its alias polynomial of degree at most $N-1$.
For use in analysis, define
@@
3.2 Choice of filter coefficients
Finally, define
@@ 3.2 Choice of filter coefficients
{pf} By Taylor’s theorem, $\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ $=y\left(\frac{(n_{i}+j)T}{n}\right)$ $=\sum_{k=0}^{d}\frac{y^{(k)}(t_{i})}{k!}\left(\frac{NT}{n}\right)^{k}\left(\frac{j}{N}\right)^{k}+R$ (8) where the remainder $R=R(i,d,n,N)$ satisfies $\left|R\right|$ $\leq R_{\text{max}}=\left(\frac{NT}{n}\right)^{d+1}\frac{M_{i}^{d}}{(d+1)!}.$ (9) We now insert this into (7). By the orthogonality property of $p^{d}_{N}$, all of the lower-order Taylor expansion terms of $\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ vanish. $\operatorname{\mathbb{E}}\hat{x}^{d}_{i}$ $=C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)\frac{x^{d}_{i}}{d!}\left(\frac{NT}{n}\right)^{d}\left(\frac{j}{N}\right)^{d}$ $\phantom{=}+C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)R(j,d,n,N)$ (10) By our choice of $C$, the first sum evaluates to $x^{d}_{i}$. $\left|\operatorname{\mathbb{E}}\hat{x}^{d}_{i}-x^{d}_{i}\right|$ $\leq CNR_{\text{max}}g_{N}^{d}$ $=\left(\frac{M_{i}^{d}T}{d+1}\right)\left(\frac{N}{n}\right)\frac{g_{N}^{d}}{h_{N}^{d}}.$
@@
{pf} Direct calculation beginning with $\begin{split}\operatorname{\operatorname{var}}\hat{x}^{d}_{j}=C^{2}\sum_{j=1}^{N}\Bigg{[}&\ \rho\left(\frac{j}{N}\right)^{2}p_{N}^{d}\left(\frac{j}{N}\right)^{2}\\ &\quad\cdot\operatorname{\operatorname{var}}w((n_{i}+j)n^{-1}T)\Bigg{]}.\end{split}$
@@
[…]
@@
Follows from applying the identity $\text{MSE}=\text{Bias}^{2}+\text{Variance}$ to Lemmas 6 and 7.
3.3 Choice of $N$
To pick the filter window size $N$, add up Thm. 8 across all $d=0,1,\ldots,m$.
@@ 3.3 Choice of $N$
4 The ridge-regression estimator
4.1 Construction of the estimator
Let $P$ be the projection matrix onto the first $m$ entries. Introduce the shorthand $\phi_{i}=\phi(Px_{i})$, $\hat{\phi}_{i}=\phi(P\hat{x}_{i})$, $\Delta\phi=\hat{\phi}_{i}-\phi_{i}$, and let $\phi$ be Hölder continuous with $\left\|\Delta\phi_{i}\right\|\leq C_{\phi}\left\|\Delta x_{i}\right\|^{\alpha}$. Let $\delta=\left\|\Delta\Phi\right\|$.
Let $\lambda\in(0,\sigma_{\text{min}}(\Phi))$ be a regularization coefficient to be determined later. Let $\Phi\in\mathbb{R}^{n^{\prime}\times p}$ be the matrix having $\phi_{i}^{\intercal},i=1,2,\ldots,n^{\prime}$ as rows, and let $\hat{\Phi}$ be the same, with $\hat{\phi}_{i}$ replacing $\phi_{i}$. Define $R,\hat{R}\in\mathbb{R}^{(p+n^{\prime})\times p}$ as the block matrices
Define $u,\hat{u}\in\mathbb{R}^{p+n^{\prime}}$ as
$u$ $=\begin{pmatrix}0_{p\times 1}\\ x^{m}_{1}\\ x^{m}_{2}\\ \vdots\\ x^{m}_{n^{\prime}}\end{pmatrix},\quad\hat{u}=\begin{pmatrix}0_{p\times 1}\\ \hat{x}^{m}_{1}\\ \hat{x}^{m}_{2}\\ \hat{\vdots}\\ \hat{x}^{m}_{n^{\prime}}\end{pmatrix}.$ (33)
$\theta$ $=R^{\dagger}u$
which motivates the estimator
$\hat{\theta}$ $=\hat{R}^{\dagger}\hat{u}.$
4.2 Error analysis
@@
{pf} By (17), $\operatorname{\mathbb{E}}\left\|v\right\|^{2}$ $=\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}^{m}\right\|^{2}$ $\leq\epsilon_{N,m}n^{\prime}n^{-\frac{2}{2m+3}}.$ (39) For all $0<r\leq 2$, Jensen’s gives $\operatorname{\mathbb{E}}\left\|v\right\|^{r}$ $\leq\frac{\epsilon_{N,m}^{r/2}(n^{\prime})^{r/2}}{n^{\frac{r}{2m+3}}}.$ (40)
@@
{pf} Bounding $\delta$ by the Frobenius norm of $\Delta\Phi$, we have $\delta^{2}$ $\leq\sum_{i=1}^{n^{\prime}}\left\|\Delta\phi_{i}\right\|^{2}$ (41) $\leq C_{\phi}^{2}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2\alpha}$ (42) By Hölder’s inequality with exponents $1/\alpha$ and $1/(1-\alpha)$, $\leq C_{\phi}^{2}\left(\sum_{i=1}^{n^{\prime}}\left(\left\|\Delta x_{i}\right\|^{2\alpha}\right)^{1/\alpha}\right)^{\alpha}\left(\sum_{i=1}^{n^{\prime}}1^{\frac{1}{1-\alpha}}\right)^{1-\alpha}$ (43) Raising both sides to the power $1/\alpha$ and taking expectations, $\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2}$ (44) $\operatorname{\mathbb{E}}\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\operatorname{\mathbb{E}}\left\|\Delta x_{i}\right\|^{2}$ (45) Applying (15), $\leq\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{n^{\frac{2}{2m+3}}}.$ (46)
To control $S$, we state and then apply a self-contained technical lemma, proven in Appendix A.
@@
[…]
@@
{pf} Applying Lemma 11 to the facts of the case, $\left\|S\right\|$ $\leq\frac{{2}\left(\delta+\lambda\right)}{\overline{\sigma}\sqrt{n^{\prime}}\left((\overline{\sigma}\sqrt{n^{\prime}}-\delta)^{+}+\lambda\right)}$ (47) As we will see that $\delta\overset{\mathbb{P}}{\longrightarrow}0$ as $n\to\infty$, the following bound is very conservative for large $n$, but it is difficult to do better when we need a high moment of $S$. $\leq\frac{{2}\left(\delta+\lambda\right)}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}$ (48) Using the fact that $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$ for $a,b\geq 0$ and $p\geq 1$, $\left\|S\right\|^{2/\alpha}$ $\leq 2^{4/\alpha-1}\frac{\delta^{2/\alpha}+\lambda^{2/\alpha}}{\lambda^{2/\alpha}\sigma_{\text{min}}^{2/\alpha}(\Phi)}$ (49) $\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}$ $\leq\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$ (50)
@@
We first state a self-contained technical lemma, proved in Appendix B, concerning Tikhonov pseudoinverses of perturbed matrices.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
{pf} We can immediately transcribe $s_{\text{head}}$ from the statement of Lemma 13. We proceed with $s_{\text{body}}$, using Chebyshev’s inequality to bound $\operatorname{\mathbb{P}}(\delta>v)$. Using $\sigma=\overline{\sigma}\sqrt{n^{\prime}}$ for temporary brevity, $3\int_{\lambda}^{\sigma}\frac{(\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v$ $\leq 3\operatorname{\mathbb{E}}\delta^{2}\int_{\lambda}^{\sigma}\frac{\lambda+\sigma}{(\lambda+\sigma-v)^{2}v^{2}\,\operatorname{d}v}$ $=\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{(\sigma+\lambda)^{2}}+\frac{\sigma-\lambda}{(\sigma+\lambda)\sigma\lambda}\right]$ “Rounding up” for simplicity, $\leq\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{\sigma^{2}}+\frac{1}{\sigma\lambda}\right]$ Applying Lemma 10 we get this lemma’s $s_{\text{body}}$. $\leq\frac{6C_{\phi}^{2}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}\sqrt{n^{\prime}}}+\frac{1}{\lambda\overline{\sigma}^{2}}\right].$ As for $s_{\text{tail}}$, by Hölder’s inequality, $\operatorname{\mathbb{E}}\frac{4\delta}{\sigma\lambda}\bm{1}_{\delta>\sigma}$ $\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\operatorname{\mathbb{P}}\left(\delta>\overline{\sigma}\sqrt{n^{\prime}}\right)\right]^{(2-\alpha)/2}$ $\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{1/\alpha}}\right]^{(2-\alpha)/2}$ $=\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{\frac{2-\alpha}{2\alpha}}}$ $=\frac{4\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$ $\leq\frac{4C_{\phi}^{2/\alpha}\epsilon_{N}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$
4.3 Bound on $\Delta\theta$

### v3

@@ A.1 Building blocks of the filter
As $\left\langle f,g\right\rangle_{N}$ converges to a Riemann integral as $N\to\infty$, so $p_{N}^{d}$ converges too. Define the filter-specific constants
@@
$g_{N}^{d}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)\left|p_{N}^{d}\left(\frac{j}{N}\right)\right|.$
Define the problem-specific constants
@@
A.2 Choice of filter coefficients
The filter is given by
@@ A.2 Choice of filter coefficients
{pf}
By Taylor’s theorem,
$\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ $=y\left(\frac{(n_{i}+j)T}{n}\right)$
$=\sum_{k=0}^{d}\frac{y^{(k)}(t_{i})}{k!}\left(\frac{NT}{n}\right)^{k}\left(\frac{j}{N}\right)^{k}+R$ (13)
where the remainder $R=R(i,d,n,N)$ satisfies
$\left|R\right|$ $\leq R_{\text{max}}=\left(\frac{NT}{n}\right)^{d+1}\frac{M_{i}^{d}}{(d+1)!}.$ (14)
We now insert this into (12). By the orthogonality property of $p^{d}_{N}$, all of the lower-order Taylor expansion terms of $\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ vanish.
$\operatorname{\mathbb{E}}\hat{x}^{d}_{i}$ $=C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)\frac{x^{d}_{i}}{d!}\left(\frac{NT}{n}\right)^{d}\left(\frac{j}{N}\right)^{d}$
$\phantom{=}+C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)R(j,d,n,N)$ (15)
By our choice of $C$, the first sum evaluates to $x^{d}_{i}$.
$\left|\operatorname{\mathbb{E}}\hat{x}^{d}_{i}-x^{d}_{i}\right|$ $\leq CNR_{\text{max}}g_{N}^{d}$
$=\left(\frac{M_{i}^{d}T}{d+1}\right)\left(\frac{N}{n}\right)\frac{g_{N}^{d}}{h_{N}^{d}}.$
@@
{pf}
Direct calculation beginning with
$\begin{split}\operatorname{\operatorname{var}}\hat{x}^{d}_{i}=C^{2}\sum_{j=1}^{N}\Bigg{[}&\ \rho\left(\frac{j}{N}\right)^{2}p_{N}^{d}\left(\frac{j}{N}\right)^{2}\\ &\quad\cdot\operatorname{\operatorname{var}}w((n_{i}+j)n^{-1}T)\Bigg{]}.\end{split}$
@@
$A_{i,N}^{d}$ $=\left(\frac{M_{i}^{d}T}{d+1}\right)^{2}\left(\frac{g_{N}^{d}}{h_{N}^{d}}\right)^{2}\quad\text{and}$
@@
Follows from applying the identity $\text{MSE}=\text{Bias}^{2}+\text{Variance}$ to Lemmas 9 and 10.
A.3 Choice of $N$
To pick the filter window size $N$, add up Thm. 11 across all $d=0,1,\ldots,m$.
@@ A.3 Choice of $N$
Appendix B Proof of Theorem 8
Let $\Delta\phi_{i}=\hat{\phi}_{i}-\phi_{i}$. By Hölder continuity, $\left\|\Delta\phi_{i}\right\|\leq C_{\phi}\left\|\Delta x_{i}\right\|^{\alpha}$. Let $\delta=\left\|\Delta\Phi\right\|$.
B.1 Error analysis
@@
{pf}
By (22),
$\operatorname{\mathbb{E}}\left\|v\right\|^{2}$ $=\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}^{m}\right\|^{2}$
$\leq\epsilon_{N,m}n^{\prime}n^{-\frac{2}{2m+3}}.$ (28)
For all $0<r\leq 2$, Jensen’s gives
$\operatorname{\mathbb{E}}\left\|v\right\|^{r}$ $\leq\frac{\epsilon_{N,m}^{r/2}(n^{\prime})^{r/2}}{n^{\frac{r}{2m+3}}}.$ (29)
@@
{pf}
Bounding $\delta$ by the Frobenius norm of $\Delta\Phi$, we have
$\delta^{2}$ $\leq\sum_{i=1}^{n^{\prime}}\left\|\Delta\phi_{i}\right\|^{2}$ (30)
$\leq C_{\phi}^{2}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2\alpha}$ (31)
By Hölder’s inequality with exponents $1/\alpha$ and $1/(1-\alpha)$,
$\leq C_{\phi}^{2}\left(\sum_{i=1}^{n^{\prime}}\left(\left\|\Delta x_{i}\right\|^{2\alpha}\right)^{1/\alpha}\right)^{\alpha}\left(\sum_{i=1}^{n^{\prime}}1^{\frac{1}{1-\alpha}}\right)^{1-\alpha}$ (32)
Raising both sides to the power $1/\alpha$ and taking expectations,
$\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2}$ (33)
$\operatorname{\mathbb{E}}\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\operatorname{\mathbb{E}}\left\|\Delta x_{i}\right\|^{2}$ (34)
Applying (20),
$\leq\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{n^{\frac{2}{2m+3}}}.$ (35)
To control $S$, we state and then apply a self-contained technical lemma, proven in Appendix C.
@@
$B$ $=\begin{pmatrix}0_{n\times n}\\ A\end{pmatrix}^{\dagger},\quad\hat{B}=\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger},\quad\text{and}$
@@
{pf}
Applying Lemma 14 to the facts of the case,
$\left\|S\right\|$ $\leq\frac{{2}\left(\delta+\lambda\right)}{\overline{\sigma}\sqrt{n^{\prime}}\left((\overline{\sigma}\sqrt{n^{\prime}}-\delta)^{+}+\lambda\right)}$ (36)
As we will see that $\delta\overset{\mathbb{P}}{\longrightarrow}0$ as $n\to\infty$, the following bound is very conservative for large $n$, but it is difficult to do better when we need a high moment of $S$.
$\leq\frac{{2}\left(\delta+\lambda\right)}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}$ (37)
Using the fact that $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$ for $a,b\geq 0$ and $p\geq 1$,
$\left\|S\right\|^{2/\alpha}$ $\leq 2^{4/\alpha-1}\frac{\delta^{2/\alpha}+\lambda^{2/\alpha}}{\lambda^{2/\alpha}\sigma_{\text{min}}^{2/\alpha}(\Phi)}$ (38)
$\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}$ $\leq\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$ (39)
@@
We first state a self-contained technical lemma, proved in Appendix D, concerning Tikhonov pseudoinverses of perturbed matrices.
@@
$B$ $=\begin{pmatrix}0_{m\times m}\\ A\end{pmatrix}^{\dagger},\quad\hat{B}=\begin{pmatrix}\lambda I_{m}\\ A+D\end{pmatrix}^{\dagger},\quad\text{and}$
@@
$\operatorname{\mathbb{E}}\left\|S\right\|$ $\leq s_{\text{head}}+s_{\text{body}}+s_{\text{tail}},$
@@
$s_{\text{head}}$ $=\frac{4\lambda}{\sigma^{2}},$
$s_{\text{body}}$ $=3\int_{\lambda}^{\sigma}\frac{(\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v,$
@@
$s_{\text{tail}}$ $=\operatorname{\mathbb{E}}\frac{4\delta}{\sigma\lambda}\bm{1}_{\delta>\sigma}.$
@@
$\operatorname{\mathbb{E}}\left\|S\right\|$ $\leq s_{\text{head}}+s_{\text{body}}+s_{\text{tail}},$
@@
$s_{\text{head}}$ $=\frac{4\lambda}{\overline{\sigma}^{2}n^{\prime}},$
$s_{\text{body}}$ $=\frac{6C_{\phi}^{2}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}\sqrt{n^{\prime}}}+\frac{1}{\lambda\overline{\sigma}^{2}}\right],$
@@
$s_{\text{tail}}$ $=\frac{4C_{\phi}^{2/\alpha}\epsilon_{N}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$
{pf}
We can immediately transcribe $s_{\text{head}}$ from the statement of Lemma 16.
We proceed with $s_{\text{body}}$, using Chebyshev’s inequality to bound $\operatorname{\mathbb{P}}(\delta>v)$. Using $\sigma=\overline{\sigma}\sqrt{n^{\prime}}$ for temporary brevity,
$3\int_{\lambda}^{\sigma}\frac{(\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v$
$\leq 3\operatorname{\mathbb{E}}\delta^{2}\int_{\lambda}^{\sigma}\frac{\lambda+\sigma}{(\lambda+\sigma-v)^{2}v^{2}\,\operatorname{d}v}$
$=\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{(\sigma+\lambda)^{2}}+\frac{\sigma-\lambda}{(\sigma+\lambda)\sigma\lambda}\right]$
“Rounding up” for simplicity,
$\leq\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{\sigma^{2}}+\frac{1}{\sigma\lambda}\right]$
Applying Lemma 13 we get this lemma’s $s_{\text{body}}$.
$\leq\frac{6C_{\phi}^{2}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}\sqrt{n^{\prime}}}+\frac{1}{\lambda\overline{\sigma}^{2}}\right].$
As for $s_{\text{tail}}$, by Hölder’s inequality,
$\operatorname{\mathbb{E}}\frac{4\delta}{\sigma\lambda}\bm{1}_{\delta>\sigma}$
$\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\operatorname{\mathbb{P}}\left(\delta>\overline{\sigma}\sqrt{n^{\prime}}\right)\right]^{(2-\alpha)/2}$
$\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{1/\alpha}}\right]^{(2-\alpha)/2}$
$=\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{\frac{2-\alpha}{2\alpha}}}$
$=\frac{4\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$
$\leq\frac{4C_{\phi}^{2/\alpha}\epsilon_{N}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$
B.2 Bound on $\Delta\theta$